At 5 percent compounded monthly, how long will it take to double your money?
Answer #1

We use the formula:
A=P(1+r/1200)^12n
where
A=future value($2x)
P=present value($x say)
r=rate of interest
n=time period.

2x=x*(1+0.05/12)^12n

(2)=(1.004167)^12n

Taking log on both sides;

log 2=12n*log (1.004167)

n=1/12[log 2/log(1.004167)]

=13.89 years(Approx).